[Hallc_running] [New Logentry] Hall C RC daily update

simona at jlab.org simona at jlab.org
Sun Aug 2 18:55:02 EDT 2020

Logentry Text:
We got tune beam in Hall C this morning. The beam profile is not adequate for running on glass cell targets: https://logbooks.jlab.org/entry/3819448 , https://logbooks.jlab.org/entry/3819430. Jay and the MCC shift crews are re-working the Hall C match. We hope to get CW for beam centering sometime during the swing shift. If beam centering is completed and if the beam is adequate, we plan to do a Moller measurement tomorrow during the day shift.

In other news the SoLID Cherenkov test went very well  (got data with tune beam): https://logbooks.jlab.org/entry/3819431 (online event display courtesy of Argonne group) , https://logbooks.jlab.org/entry/3819462 , https://logbooks.jlab.org/entry/3819492. Data taking will continue with CW when available. 

We had a RC meeting today at 4:15 pm. Thanks to those who connected.


This is a plain text email for clients that cannot display HTML.  The full logentry can be found online at https://logbooks.jlab.org/entry/3819574
-------------- next part --------------
An HTML attachment was scrubbed...
URL: <https://mailman.jlab.org/pipermail/hallc_running/attachments/20200802/8bc6bd61/attachment.html>

More information about the Hallc_running mailing list